# Endmember Extraction from Hyperspectral Images Using Self-Dictionary Approach with Linear Programming

### v2

@@ Abstract
Hyperspectral imaging, endmember extraction, nonnegative matrix factorization, linear mixing model, pure pixel, self-dictionary, linear programming, column generation method.
@@ I Introduction
Many zero elements may exist in the optimal solution of a Hottopixx model. This suggests that Hottopixx models can be solved by breaking them up into smaller subproblems and solving them. We present Theorem 2, which shows that the optimal solution of the full Hottopixx model can be obtained by solving its subproblems. We then develop an RCE algorithm based on this theoretical result. The details are described in Section IV.
@@
The postprocessing step in the Hottopixx method described in [28] conducts data clustering and outputs one element from each cluster. The choice of elements affects the endmember extraction performance of Hottopixx. The methods used in the previous studies [16, 19, 28] were based on the optimal solutions to Hottopixx models. Alternatively, we propose a method based on the shape of clusters; it computes the centroids of each cluster and chooses elements close to them. This method is simple but effective in enhancing the endmember extraction performance of Hottopixx. We will call it the cluster centroid choice. The details are described in Section V.
@@
We developed EEHT by incorporating RCE and the cluster centroid choice. We then experimentally tested its performance on endmember extraction problems. The first experiment (Section VI-A) examined the computational time of RCE. The results show that RCE is faster than the direct method of solving Hottopixx models. The second experiment (Section VI-B) evaluated the endmember extraction performance of EEHT on semi-real HSI datasets. We compared EEHT with six existing methods, including MERIT from [30], which is a method based on a self-dictionary, similar to EEHT. The results show that, in terms of the estimation accuracy of the endmember signatures, the performance of EEHT is comparable to MERIT when the penalty parameter of MERIT is properly adjusted. Finally, we conducted an experimental study (Section VI-C) on hyperspectral unmixing of the Urban HSI dataset. The results show that, compared with the existing methods, EEHT usually provides more accurate estimations of endmember signatures for Urban.
@@ I-A Organization of This Paper
This paper is organized as follows. Section II is devoted to the preliminaries. Section III formulates the endmember extraction problem and explains the algorithm of Hottopixx. It also reviews related methods. Section IV and Section V present the details of the RCE and EEHT. Section VI describes the experiments.
II Preliminaries
II-A Notation
$\bm{0}$ denotes a vector of all zeros, $\bm{1}$ a vector of all ones, $O$ a matrix of all zeros, $I$ the identity matrix, and $J$ a matrix of all ones. +$\mathbb{R}^{d\times n}_{+}$ represents the set of nonnegative matrices.
@@ II-A Notation
We use $\mathcal{N}$ to denote $\{1,\ldots,n\}$. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$. For $\mathcal{I}\subset\mathcal{N}$, we use $A(\mathcal{I})$ or $[\bm{a}_{i}:i\in\mathcal{I}]$ to denote the submatrix of $A$ indexed by $\mathcal{I}$ in ascending order. Specifically, if =$\mathcal{I}=\{i_{1},\ldots,i_{p}\}$ with $i_{1}<\cdots<i_{p}$, then $A(\mathcal{I})=[\bm{a}_{i}:i\in\mathcal{I}]=[\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{p}}]$.
[…]
@@
II-B Duality of LP Problems
As mentioned in Section I, we propose the RCE algorithm for solving Hottopixx models efficiently. This algorithm is founded on the duality of LP problems. For $A\in\mathbb{R}^{d\times n},\bm{b}\in\mathbb{R}^{d}$ and $\bm{c}\in\mathbb{R}^{n}$, consider the primal and dual pair of LP,
(Primal) $\min_{\bm{x}\in\mathbb{R}^{n}}$ $\bm{c}^{\top}\bm{x}$ s.t. $A\bm{x}=\bm{b},\ \bm{x}\geq\bm{0}.$
(Dual) $\max_{\bm{y}\in\mathbb{R}^{n}}$ $\bm{b}^{\top}\bm{y}$ s.t. $A^{\top}\bm{y}\leq\bm{c}.$
  • The duality theorem: If the primal and dual are feasible, then, there exist optimal solutions $\bm{x}$ and $\bm{y}$ to the primal and dual and they satisfy $\bm{c}^{\top}\bm{x}=\bm{b}^{\top}\bm{y}$.
  • Implication of the weak duality theorem: Let $\bm{x}$ and $\bm{y}$ be feasible solutions to the primal and dual. If $\bm{c}^{\top}\bm{x}=\bm{b}^{\top}\bm{y}$, then, $\bm{x}$ and $\bm{y}$ are optimal solutions to the primal and dual.
For more details, refer to the textbooks [4, 3]. We will use these results to develop RCE.
III Self-Dictionary Approaches for Solving Endmember Extraction Problems
III-A Endmember Extraction of HSIs
Suppose that we are given an HSI data for some area consisting of $n$ pixels acquired by a hyperspectral camera with $d$ spectral bands. We represent the HSI as a matrix $A\in\mathbb{R}^{d\times n}$ where the $(i,j)$th entry of $A$ stores the measurement at the $i$th band and $j$th pixel. In this paper, we refer to such a matrix $A$ an HSI matrix. The $j$th column $\bm{a}_{j}$ of $A$ corresponds to the observed spectral signature at the $j$th pixel.
Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$ be an HSI matrix and let us consider a linear mixing model (LMM) for $A$. Under the LMM, the observed spectral signatures $\bm{a}_{1},\ldots,\bm{a}_{n}$ at each pixel are written as
@@ III-A Endmember Extraction of HSIs
The pure pixel assumption is considered reasonable. Meanwhile, it is difficult to determine the exact number of endmembers in an HSI. Thus, in our experiments we estimate it in the way shown in previous studies. Problem 1 is known to be equivalent to the problem of computing a nonnegative matrix factorization (NMF) under the separability assumption. Arora et al. [2] studied the separable NMF problems from a theoretical point of view. Their results tell us that we can find $W$ from $A$ exactly if $A$ is noiseless; otherwise, if $A$ contains noise $V$ so that the amount of $V$ is smaller than some level, we can find $r$ columns of $A$ close to those of $W$. Given $A$ as shown in (3), we say that an algorithm for Problem 1 is robust to noise if it can find $r$ such columns of $A$. For further reading on separable NMF, refer to [12, 18].
@@ III-B Hottopixx Methods
In this paper, we call $\mathsf{H}$ the Hottopixx model, although it differs from the original Hottopixx model proposed by Bittorf et al. [6]. As shown in Section IV-B, $\mathsf{H}$ can be reduced to an LP problem. The previous study [28] directly solved this LP using a CPLEX solver. The constraints of $\mathsf{H}$ relax the constraint $\|X\|_{\mathrm{row},0}=r$ of problem (5). However, it is possible for many rows of the optimal solution $X$ in $\mathsf{H}$ to be zero, since the constraints of $\mathsf{H}$ imply that the $i$th row of $X$ is zero if $X(i,i)=0$ and many of $X(1,1),\ldots,X(n,n)$ could be zero.
@@
The second step involves choosing $r$ columns of $A$ using the optimal solution $X$ of $\mathsf{H}$. A simple way of doing so is to find a set $\mathcal{I}$ of $r$ indices corresponding to the $r$ largest elements of $\mathrm{diag}(X)$ and, then, output columns $\bm{a}_{i}$ with $i\in\mathcal{I}$. Theorem 3.1 of [28] ensures that the algorithm is robust to noise. However, the theorem is invalid and does not hold in a case where there are overlaps of pure pixels in the HSIs. To cope with this issue, we use a clustering technique. Algorithm 1 describes each step of Hottopixx with postprocessing [28]. A cluster $\mathcal{S}_{\ell}$ in the algorithm refers to a subset of the column index set $\mathcal{N}$ of $A$. Due to page limitations, a detailed description of how to construct clusters is provided in Appendix B.
@@
      Construct a cluster $\mathcal{S}_{\ell}$ using $\bm{p}_{\ell}$; see Appendix B for the details.
@@
Algorithm 1 has an advantage over other algorithms [6, 16, 19] in that it does not require us to specify the noise level involved in $A$ as input. Theorem 3.2 of [28] ensures that Algorithm 1 is robust to noise. Even if there are overlaps among pure pixels in the HSI, the theorem remains valid and the algorithm is robust to noise. In most cases, there are a number of pixels in the HSIs that are close to pure pixels. Specifically, we will see this feature in the Urban dataset in Section VI-C. Hence, when solving Problem 1, we should use the postprocessing procedure described in step 2 of Algorithm 1.
@@
III-C Related Methods
So far, many methods have been proposed for addressing Problem 1, as detailed in the survey papers [5, 26] and the textbook [18]. These methods can be categorized into two groups: convex optimization-based methods and greedy methods. Hereinafter, we will use the term “convex method” as a shorthand to refer to the convex optimization-based method. Convex methods are robust to noise but computationally expensive, whereas greedy methods are faster but less robust. Hottopixx is an example of a convex method.
III-C1 Convex Methods
As mentioned earlier, Problem 1 can be formulated as a sparse optimization problem. However, directly solving sparse optimization problems is computationally intractable. Thus, we instead solve their convex relaxation problems. A typical algorithm for convex methods involves computing the optimal solution to the convex relaxation problem and subsequently extracting $r$ columns of $A$ using this optimal solution.
Our approach is closely related to FGNSR, proposed by Gillis and Luce [20], and MERIT, proposed by Nguyen, Fu and Wu [30], both of which are convex methods using self-dictionaries. Here, let us recall problem (5), which is the sparse optimization formulation for Problem 1. FGNSR and MERIT solve its convex relaxation as follows:
@@ III-C1 Convex Methods
for a penalty parameter $\lambda\geq 0$, where $\Phi$ is the regularization term and $\mathcal{G}$ is the feasible region. Here, $\Phi$ is designed to minimize the number of nonzero rows in $X$.
@@
where $c_{i}=\|A(:,i)\|_{1}$. Problem (6) with $\Phi$ and $\mathcal{G}$ as described above is closely related to the original Hottopixx model by Bittorf et al. [6]. Gillis and Luce [20] showed that the projection of a matrix onto $\mathcal{G}$ can be computed in $O(n^{2}\log n)$. Taking that into account it, they developed a first-order method equipped with Nesterov’s acceleration to solve these problems.
@@
Convex methods for solving Problem 1 and related problems can be found in the papers by Elhamifar, Sapiro and Vidal [10] and by Esser et al. [11], both published in 2012. Fu and Ma [13] analyzed the performance of convex method based on problem (6) with $\Phi(X)=\|X\|_{\infty,q}$ for $0<q\leq 1$ and with $\mathcal{G}$ defined by the simplex constraint. Fu, Sidiropoulos and Ma [14] applied convex optimization approaches to power spectral separation.
III-C2 Greedy Methods
Greedy methods choose the columns of the HSI matrix one by one on the basis of certain criteria. One of the most popular greedy methods for solving Problem 1 is SPA, which dates back to the work of Araújo et al. [1] in 2001. Subsequently, Gillis and Vavasis [21] gave a theoretical justification for the robustness of SPA to noise. SPA recursively chooses the columns of the input matrix. For a column chosen in the previous iteration, it projects columns onto the orthogonal complement of the chosen column and then chooses one with the maximum $L_{2}$ norm. To enhance the robustness of SPA to noise, Gillis and Vavasis [22] developed PSPA, which is SPA with preconditioning, and Mizutani [27] ER, which is SPA with a preprocessing. Both preconditioning and preprocessing use singular value decomposition (SVD) based dimensionality reduction and minimum volume ellipsoids. Gillis [17] proposed SNPA, which is similar to that of SPA but it uses a projection onto the convex hull of the origin and previously chosen columns. Nascimento and Bioucas-Dias [29] proposed VCA, which applies an SVD-based dimensionality reduction to the input matrix, and then, runs the same algorithm as SPA except for the criterion for choosing columns. Other popular methods in the remote sensing community include PPI [7], N-FINDR [31], and AVMAX and SVMAX [9].
IV Solving Hottopixx Models Efficiently
IV-A Algorithm Outline
In Section IV-B, we show that $\mathsf{H}$ can be reduced to an LP with $O(n^{2})$ variables and $O(n^{2})$ constraints for HSI matrices $A\in\mathbb{R}^{d\times n}$ with $d\leq n$. Accordingly, when $n$ is large, directly solving the LP, as in the previous study [28], becomes computationally challenging. To remedy this computational issue, we develop the RCE algorithm for solving $\mathsf{H}$ efficiently based on the framework of column generation, which is known as a classical but powerful technique for solving large-scale LPs; see, for instance, the textbook [4] for details on column generation methods. The algorithm is based on the observation that there may be many zero rows in the optimal solution of $\mathsf{H}$. We can significantly reduce the size of $\mathsf{H}$ by exploiting the sparsity of the optimal solution.
Let $\mathcal{L}$ and $\mathcal{M}$ be the subsets of $\mathcal{N}$. Hereinafter, we will use $\ell$ and $m$ to denote $|\mathcal{L}|$ and $|\mathcal{M}|$, respectively. RCE uses a subproblem of $\mathsf{H}$ with a variable $X\in\mathbb{R}^{\ell\times m}$. The feasible region $\mathcal{F}(\ell,m)$ is defined by

### v3

@@ Abstract
Hyperspectral imaging, endmember extraction, nonnegative matrix factorization, linear mixing model, bilinear mixing model, pure pixel, self-dictionary, linear programming, column generation method.
@@ I Introduction
Many zero elements may exist in the optimal solution of a Hottopixx model. This suggests that Hottopixx models can be solved by breaking them up into smaller subproblems and solving them. We present Theorem 2, which shows that the optimal solution of the full Hottopixx model can be obtained by solving its subproblems. We then develop an RCE algorithm based on this theoretical result. The details are described in Section V.
@@
The postprocessing step in the Hottopixx method described in [31] conducts data clustering and outputs one element from each cluster. The choice of elements affects the endmember extraction performance of Hottopixx. The methods used in the previous studies [17, 20, 31] were based on the optimal solutions to Hottopixx models. Alternatively, we propose a method based on the shape of clusters; it computes the centroids of each cluster and chooses elements close to them. This method is simple but effective in enhancing the endmember extraction performance of Hottopixx. We will call it the cluster centroid choice. The details are described in Section VI.
@@
We developed EEHT by incorporating RCE and the cluster centroid choice. We then experimentally tested its performance on endmember extraction problems. The first experiment (Section VII-A) examined the computational time of RCE. The second experiment (Section VII-B) evaluated the endmember extraction performance of EEHT on synthetic HSI datasets. These datasets were constructed from two real HSIs, Jasper Ridge and Samson. We compared EEHT with six existing methods, including MERIT from [33], which is a method based on a self-dictionary, similar to EEHT. Finally, we conducted an experimental study (Section VII-C) on hyperspectral unmixing of the Urban HSI dataset. The results show that, compared with the existing methods, EEHT usually provides more accurate estimates of endmember signatures for Urban.
@@ I-A Organization of This Paper
This paper is organized as follows. Section II reviews related work on endmember extraction methods. Section III introduces the preliminaries. Section IV formulates the endmember extraction problem and explains the algorithm of Hottopixx. Section V and Section VI present the details of the RCE and EEHT. Section VII describes the experiments.
II Related Work
Numerous methods have been proposed for extracting endmembers from HSIs, as detailed in the survey papers [5, 29] and the textbook [19]. These methods can be categorized into two groups: convex optimization-based methods and greedy methods. Hereinafter, we use the term “convex method” as a shorthand to refer to the convex optimization-based method. Convex methods are robust to noise but computationally expensive, whereas greedy methods are faster but less robust. Hottopixx is an example of a convex method.
II-1 Convex Methods
As discussed in Section IV-B, the endmember extraction problem can be formulated as a sparse optimization problem. However, directly solving sparse optimization problems is computationally intractable. Thus, we instead solve their convex relaxation problems. A typical convex method first computes the optimal solution to the relaxed problem and then identifies the endmembers based on this solution. Hottopixx is closely related to FGNSR, proposed by Gillis and Luce [21], and MERIT, proposed by Nguyen, Fu and Wu [33], both of which are convex methods using self-dictionaries. We will review FGNSR and MERIT in more detail in Section IV-C.
II-2 Greedy Methods
Greedy methods choose the columns of the HSI matrix one by one on the basis of certain criteria. One of the most popular greedy methods for endmember extraction is SPA, which dates back to the work of Araújo et al. [1] in 2001. Subsequently, Gillis and Vavasis [22] gave a theoretical justification for the robustness of SPA to noise. SPA recursively chooses columns of the input matrix. For a column chosen in the previous iteration, it projects columns onto the orthogonal complement of the chosen column and then chooses one with the maximum $L_{2}$ norm. To enhance the robustness of SPA to noise, Gillis and Vavasis [23] developed PSPA, which is SPA with preconditioning, and Mizutani [30] ER, which is SPA with a preprocessing. Both preconditioning and preprocessing use singular value decomposition (SVD) based dimensionality reduction and minimum volume ellipsoids. Gillis [18] proposed SNPA, which is similar to SPA but it uses a projection onto the convex hull of the origin and previously chosen columns. Nascimento and Bioucas-Dias [32] proposed VCA, which applies an SVD-based dimensionality reduction to the input matrix, and then, runs the same algorithm as SPA except for the criterion for choosing columns. Other popular methods in the remote sensing community include PPI [7], N-FINDR [34], and AVMAX and SVMAX [9].
III Preliminaries
III-A Notation
$\bm{0}$ denotes a vector of all zeros, $\bm{1}$ a vector of all ones, $O$ a matrix of all zeros, $I$ the identity matrix, and $J$ a matrix of all ones. $\mathbb{R}^{d\times n}_{+}$ represents the set of nonnegative matrices.
@@ III-A Notation
We use $\mathcal{N}$ to denote $\{1,\ldots,n\}$. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$. For $\mathcal{I}\subset\mathcal{N}$, we use $A(\mathcal{I})$ or $\bm{a}_{i}\mid i\in\mathcal{I}]$ to denote the submatrix of $A$ indexed by $\mathcal{I}$ in ascending order. Specifically, if $\mathcal{I}=\{i_{1},\ldots,i_{p}\}$ with $i_{1}<\cdots<i_{p}$, then $A(\mathcal{I})=[\bm{a}_{i}\mid i\in\mathcal{I}]=[\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{p}}]$.
We use $\|\cdot\|$ to denote any norm of a vector or a matrix. In particular, $\|\cdot\|_{p}$ refers to the $L_{p}$-norm of a vector or a matrix, and $\|\cdot\|_{F}$ the Frobenius norm of a matrix. For $A\in\mathbb{R}^{d\times n}$, we define $\|A\|_{p,q}=\sum_{i=1}^{n}\|A(i,:)\|_{p}^{q}$ and use $\|A\|_{\mathrm{row},0}$ to denote the number of nonzero rows in $A$. Moreover, $\langle A,B\rangle$ indicates the inner product of two matrices $A$ and $B$ of the same size, defined by $\langle A,B\rangle=\mathrm{tr}(A^{\top}B)$.
@@
III-B Duality and Solution Methods for LP Problems
As mentioned in Section I, we propose the RCE algorithm for solving Hottopixx models efficiently. This algorithm is based on the duality of LP problems. For $A\in\mathbb{R}^{m\times n}$, $\bm{b}\in\mathbb{R}^{m}$, and $\bm{c}\in\mathbb{R}^{n}$, consider an LP problem in the standard form and its corresponding dual problem, given as follows:
(Primal) $\min_{\bm{x}\in\mathbb{R}^{n}}$ $\bm{c}^{\top}\bm{x}$ s.t. $A\bm{x}=\bm{b},\ \bm{x}\geq\bm{0},$
(Dual) $\max_{\bm{y}\in\mathbb{R}^{m}}$ $\bm{b}^{\top}\bm{y}$ s.t. $A^{\top}\bm{y}\leq\bm{c}.$
  • The duality theorem: If the primal and dual are feasible, then there exist optimal solutions $\bm{x}$ and $\bm{y}$ to the primal and dual and they satisfy $\bm{c}^{\top}\bm{x}=\bm{b}^{\top}\bm{y}$.
  • Implication of the weak duality theorem: Let $\bm{x}$ and $\bm{y}$ be feasible solutions to the primal and dual. If $\bm{c}^{\top}\bm{x}=\bm{b}^{\top}\bm{y}$, then $\bm{x}$ and $\bm{y}$ are optimal solutions to the primal and dual.
For more details, see the textbooks [4, 3]. We will use these results to develop RCE.
In solving LPs, the simplex method and the interior-point method are mainly used. The iteration complexity has been established for the interior-point method. For instance, the primal-dual interior-point method computes the optimal solutions of the primal and dual problems simultaneously, where the iteration count is $O(\sqrt{n}\log(1/\epsilon))$ for a given tolerance $\epsilon>0$, and each iteration involves solving a system of linear equations with $2n+m$ variables and $2n+m$ equations; see the textbook [35] for further details.
IV Self-Dictionary Approaches for Solving Endmember Extraction Problems
IV-A Endmember Extraction of HSIs
Suppose that we are given HSI data for some area consisting of $n$ pixels acquired by a hyperspectral camera with $d$ spectral bands. We represent the HSI as a matrix $A\in\mathbb{R}^{d\times n}$ where the $(i,j)$th entry of $A$ stores the measurement at the $i$th band and $j$th pixel. In this paper, we refer to such a matrix $A$ an HSI matrix. The $j$th column $\bm{a}_{j}$ of $A$ corresponds to the observed spectral signature at the $j$th pixel.
Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$ be an HSI matrix, and consider a linear mixing model (LMM) for $A$. Under the LMM, the observed spectral signatures $\bm{a}_{1},\ldots,\bm{a}_{n}$ at each pixel are written as
@@ IV-A Endmember Extraction of HSIs
The pure pixel assumption is considered reasonable. Meanwhile, it is difficult to determine the exact number of endmembers in an HSI. Thus, in our experiments, we use the estimate of the number of endmembers provided by previous studies. Problem 1 is known to be equivalent to the problem of computing a nonnegative matrix factorization (NMF) under the separability assumption. Arora et al. [2] studied the separable NMF problems from a theoretical point of view. Their results tell us that we can find $W$ from $A$ exactly if $A$ is noiseless; otherwise, if $A$ contains noise $V$ so that the amount of $V$ is smaller than some level, we can find $r$ columns of $A$ close to those of $W$. Given $A$ as shown in (3), we say that an algorithm for Problem 1 is robust to noise if it can find $r$ such columns of $A$. For further reading on separable NMF, refer to [13, 19].
@@ IV-B Hottopixx Methods
In this paper, we call $\mathsf{H}$ the Hottopixx model, although it differs from the original Hottopixx model proposed by Bittorf et al. [6]. As shown in Section V-B, $\mathsf{H}$ can be reduced to an LP problem. The previous study [31] directly solved this LP using a CPLEX solver. The constraints of $\mathsf{H}$ relax the constraint $\|X\|_{\mathrm{row},0}=r$ of problem (5). However, it is possible for many rows of the optimal solution $X$ in $\mathsf{H}$ to be zero, since the constraints of $\mathsf{H}$ imply that the $i$th row of $X$ is zero if $X(i,i)=0$ and many of $X(1,1),\ldots,X(n,n)$ could be zero.
@@
The second step involves choosing $r$ columns of $A$ using the optimal solution $X$ of $\mathsf{H}$. A simple way of doing so is to find a set $\mathcal{I}$ of $r$ indices corresponding to the $r$ largest elements of $\mathrm{diag}(X)$ and, then, output columns $\bm{a}_{i}$ with $i\in\mathcal{I}$. Theorem 3.1 of [31] ensures that the algorithm is robust to noise. However, the theorem is invalid and does not hold in a case where there are overlaps of pure pixels in the HSIs. To cope with this issue, we use a clustering technique. Algorithm 1 describes each step of Hottopixx with postprocessing [31]. A cluster $\mathcal{S}_{\ell}$ in the algorithm refers to a subset of the column index set $\mathcal{N}$ of $A$. Due to page limitations, a detailed description of how to construct clusters is provided in Appendix D.
@@
      Construct a cluster $\mathcal{S}_{\ell}$ using $\bm{p}_{\ell}$; see Appendix D for the details.
@@
Algorithm 1 has an advantage over other algorithms [6, 17, 20] in that it does not require us to specify the noise level involved in $A$ as input. Theorem 3.2 of [31] ensures that Algorithm 1 is robust to noise. Even if there are overlaps among pure pixels in the HSI, the theorem remains valid and the algorithm is robust to noise. In most cases, there are a number of pixels in the HSIs that are close to pure pixels. Specifically, we will see this feature in the Urban dataset in Section VII-C. Hence, when solving Problem 1, we should use the postprocessing procedure described in step 2 of Algorithm 1.
@@
IV-C Convex Methods in Prior Work
We review convex methods from prior work, focusing in particular on FGNSR [21], and MERIT [33]. To begin, we recall problem (5), which is the sparse optimization formulation for Problem 1. FGNSR and MERIT solve its convex relaxation as follows:
@@ IV-C Convex Methods in Prior Work
for a penalty parameter $\lambda\geq 0$, where $\Phi$ is the regularization term and $\mathcal{G}$ is the feasible region. Here, $\Phi$ is designed to promote row sparsity in $X$, that is, to minimize the number of nonzero rows.
@@
where $c_{i}=\|A(:,i)\|_{1}$. Problem (6) with $\Phi$ and $\mathcal{G}$ as described above is closely related to the original Hottopixx model by Bittorf et al. [6]. Gillis and Luce [21] showed that the projection of a matrix onto $\mathcal{G}$ can be computed in $O(n^{2}\log n)$. Taking this into account it, they developed a first-order method equipped with Nesterov’s acceleration to solve these problems.
@@
Convex methods for solving Problem 1 and related problems can be found in the works of Elhamifar, Sapiro and Vidal [11], and of Esser et al. [12]. Fu and Ma [14] analyzed the performance of convex method based on problem (6) with $\Phi(X)=\|X\|_{\infty,q}$ for $0<q\leq 1$ and with $\mathcal{G}$ defined by the simplex constraint. Fu, Sidiropoulos and Ma [15] applied convex optimization approaches to power spectral separation.
V Solving Hottopixx Models Efficiently
V-A Algorithm Outline
In Section V-B, we show that $\mathsf{H}$ can be reduced to an LP with $O(n^{2})$ variables and $O(n^{2})$ constraints for HSI matrices $A\in\mathbb{R}^{d\times n}$ with $d\leq n$. Accordingly, when $n$ is large, directly solving the LP, as in the previous study [31], becomes computationally challenging. To remedy this computational issue, we develop the RCE algorithm for solving $\mathsf{H}$ efficiently based on the framework of column generation, which is known as a classical but powerful technique for solving large-scale LPs; see, for instance, the textbook [4] for details on column generation methods. The algorithm is based on the observation that there may be many zero rows in the optimal solution of $\mathsf{H}$. We can significantly reduce the size of $\mathsf{H}$ by exploiting the sparsity of the optimal solution.
Let $\mathcal{L}$ and $\mathcal{M}$ be the subsets of $\mathcal{N}$, where recall that $\mathcal{N}$ is defined as $\{1,\ldots,n\}$ in Section III-A. Hereinafter, we use $\ell$ and $m$ to denote $|\mathcal{L}|$ and $|\mathcal{M}|$, respectively. RCE uses a subproblem of $\mathsf{H}$ with a variable $X\in\mathbb{R}^{\ell\times m}$. The feasible region $\mathcal{F}(\ell,m)$ is defined by